Verify rank nullity theorem for the linear 

transformation T : R

3 → R

3

 defined by : 9 

T(x, y, z) = (x + 2y – z, y + z, x + y – 2z)


1
Expert's answer
2021-06-24T11:51:40-0400

"T:\mathbb{R}^3\rightarrow \mathbb{R}^3" , "T(x,y,z)=(x+2y-z,\ y+z,\ x+y-2z)"

Let us find the dimension of the image:

"T(x,y,z)=\begin{pmatrix}
 1 & 2&-1 \\
 0 & 1&1\\
1&1&-2
\end{pmatrix}\begin{pmatrix}
 x \\
 y \\
z
\end{pmatrix}=x\begin{pmatrix}
 1\\0\\1
\end{pmatrix}+ y\begin{pmatrix}
 2\\1\\1
\end{pmatrix}+z\begin{pmatrix}
 -1\\1\\-2
\end{pmatrix}"

"\text{Im} \ T=\text{Span}\begin{Bmatrix}
 \begin{pmatrix}
 1\\0\\1
\end{pmatrix}, \begin{pmatrix}
 2\\1\\1
\end{pmatrix}, \begin{pmatrix}
 - 1\\1\\-2
\end{pmatrix}
\end{Bmatrix}= \text{Span}\begin{Bmatrix}
 \begin{pmatrix}
 1\\0\\1
\end{pmatrix}, \begin{pmatrix}
 2\\1\\1
\end{pmatrix}, \begin{pmatrix}
 2\\1\\1
\end{pmatrix}-3 \begin{pmatrix}
 1 \\0\\
 1
\end{pmatrix}
\end{Bmatrix}= \text{Span}\begin{Bmatrix}
 \begin{pmatrix}
 1\\0\\1
\end{pmatrix},
\begin{pmatrix}2\\1\\1 \end{pmatrix}
\end{Bmatrix}"

"\text{dim } \text{Im} \ T=2"


Now we find the dimension of the kernel:

"T(x,y,z)=(x+2y-z,\ y+z,\ x+y-2z)=(0,0,0)"

We have that "y=-z, \ \ x=z-2y=3z" .

So, "\text{Ker} \ T=\text{Span} \begin{Bmatrix}
\begin{pmatrix}
3\\-1\\1\end{pmatrix}
\end{Bmatrix}"

"\text{dim Ker}\ T=1"


The rank-nullity theorem:

The dimension of the kernel plus the dimension of the image equals the dimension of the domain.

In this case we have "\text{dim Ker} \ T+\text{dim Im}\ T=\text{dim} \ \mathbb{R}^3" and "1+2=3" .